Krell Industries has a share price of today. If Krell is expected to pay a dividend of

this year, and its stock price is expected to grow to at the end of the year, what is Krell's dividend yield and equity cost of capital?


What is Krell's dividend yield?

What is Krell's capital gain rate?

What is Krell's equity cost of capital?
Answer #1

dividend yield: 0.67/21.87= 0.03063557738 or answer: 3.1%


capital gain: (24.35-21.87)/21.87= 0.113397348 or answer: 11.3%


cost of capital: 3.1+ 11.3= answer: 14.4%

answered by: Andrew San Andres